### v2

# BPO: Towards Balanced Preference Optimization between Knowledge Breadth and Depth in Alignment

Abstract
Reinforcement Learning with Human Feedback (RLHF) is the key to the success of large language models (LLMs) in recent years. In this work, we first introduce the concepts of knowledge breadth and knowledge depth, which measure the comprehensiveness and depth of an LLM or knowledge source, respectively. We reveal that the imbalance in the number of instructions and responses can lead to a potential disparity in breadth and depth learning within alignment tuning datasets by showing that even a simple uniform method for balancing the number of instructions and responses can lead to significant improvements. Building on this, we further propose Balanced Preference Optimization (BPO), designed to dynamically augment the knowledge depth of each sample. BPO is motivated by the observation that the usefulness of knowledge varies across samples, necessitating tailored learning of knowledge depth. To achieve this, we introduce gradient-based clustering, estimating the knowledge informativeness and usefulness of each augmented sample based on the model’s optimization direction. Our experimental results on various benchmarks demonstrate that BPO outperforms other baseline methods in alignment tuning while maintaining training efficiency. Furthermore, we conduct a detailed analysis of each component of BPO, providing guidelines for future research in preference data optimization. Our code will be published soon on https://github.com/Sizhe04/Balanced-Preference-Optimization.
BPO: Towards Balanced Preference Optimization between Knowledge Breadth and Depth in Alignment
Sizhe Wang1  Yongqi Tong2  Hengyuan Zhang3  Dawei Li4  Xin Zhang2  Tianlong Chen5 1University of Southern California, 2Ant Group, 3Tsinghua University 4Arizona State Univeristy, 5University of North Carolina at Chapel Hill sizhewan@usc.edu, {tongyongqi.yq, evan.zx}@antgroup.com, zhang-hy22@mails.tsinghua.edu.cn, daweili5@asu.edu, tianlong@cs.unc.edu
1 Introduction
Figure 1: Overview of knowledge breadth and depth, and how we link them with the number of instructions and responses in alignment tuning datasets.
Reinforcement Learning with Human Feedback (RLHF) Christiano et al. (2017) has played a pivotal role in the success of large language models (LLMs) in recent years. It aims to align LLMs with human values and preferences during the post-training phase, leveraging extensive pairwise feedback from human annotators. To boost this alignment tuning, various advanced training techniques have been introduced, including Direct Preference Optimization (DPO) Rafailov et al. (2024), Sequence Likelihood Calibration (SLiC) Zhao et al. (2022), and Kahneman-Tversky Optimization (KTO) Ethayarajh et al. (2024).
Another line of effort aims to enhance the alignment process from a data perspective. This involves techniques such as data selection Zhou et al. (2024), sampling Khaki et al. (2024a), and rearrangement Pattnaik et al. (2024a) to enhance various aspects of the original preference learning dataset, including quality Zhou et al. (2024), diversity Song et al. (2024a), and difficulty Liu et al. (2023); Pattnaik et al. (2024a). While existing data optimization methods focus on individual components of preference learning data (either the instruction or the response), a unified approach to systematically optimize preference data as a whole remains lacking.
To address this problem, we begin by proposing the concepts of knowledge breadth and knowledge depth, which measure the comprehensiveness and depth of an LLM or knowledge source respectively. Building on these concepts, we conduct a systematic analysis of the resources allocated for enhancing the breadth and depth of knowledge during alignment tuning accordingly, uncovering a potential imbalance stemming from the structure of alignment datasets. To validate this imbalance, we propose a simple balance method, that includes knowledge breadth compression (KBC) and knowledge depth augmentation (KDA), achieved by uniformly reducing the number of prompts and increasing the number of response pairs in the preference learning dataset. Figure 1 demonstrates the overview process. The substantial improvement brought by both KBC and KDA confirms the effectiveness of our balance approach in alignment tuning datasets.
We further examine the preference learning dataset and observe the need to dynamically augment the knowledge depth of different samples. Some samples in preference data contain rich, informative knowledge, while others consist of relatively shallow or less useful information. The former deserves more response pairs to allow LLMs to fully learn from them, whereas prior studies Zhou et al. (2024) have shown that excessive focus on less useful knowledge can lead to negative outcomes. Therefore, we propose Balanced Preference Optimization (BPO), a method designed to dynamically augment the knowledge depth of each sample based on its informativeness. BPO first follows a general pipeline Albalak et al. (2024); Xia et al. (2024) to sample a subset of representative and diverse prompts by clustering. Next, we introduce gradient-based Pruthi et al. (2020); Han et al. (2023) clustering to estimate the optimal knowledge depth for each sample from the perspective of model optimization. By clustering each augmented (prompt-response pair) based on their gradient features and selecting those closest to the cluster centers, BPO ensures that samples with gradient features near the centroids receive more focus. These centrally positioned samples are typically more informative and helpful for model optimization, making them deserving of greater knowledge-depth learning resources.
Extensive experiments across various benchmarks demonstrate both the effectiveness and computational efficiency of our BPO approach compared to other baseline methods. Additional ablation studies and hyperparameter analyses confirm the effectiveness of each component, highlighting BPO’s robustness across different hyperparameter settings. We also investigate alternative methods for dynamically estimating the required knowledge depth, with experimental results showing that the gradient-based approach performs best, achieving the highest human preference scores.
To summarize, our contribution in this work is threefold:
  • We introduce the concept of balancing knowledge breadth and depth, and conduct preliminary experiments to validate the effectiveness of this balancing in alignment tuning datasets.
  • We propose BPO, which involves estimating samples’ difficulties and dynamically augmenting the knowledge depth for each sample via hierarchical sampling.
  • Through extensive experiments, we show that BPO outperforms previous preference data optimization methods. We also conduct further exploration on BPO to provide in-depth hints.
Figure 2: Preliminary experiment results on SafeRLHF using simple balance.
2 Preliminary
In this section, we first provide the definition of knowledge breadth and depth and bridge the connection between the two concepts with instruction and response numbers in alignment tuning datasets. Then, we propose a simple strategy to balance the knowledge breadth and depth in the DPO tuning process and conduct preliminary experiments to evaluate its effectiveness.
2.1 Knowledge Breadth and Depth
In general, knowledge breadth $\mathcal{B}$ represents the model’s range of knowledge across various subjects or domains, which is a measure of how many different areas or topics the model can understand and provide information on. In contrast, knowledge depth $\mathcal{K}$ refers to how well an LLM can provide in-depth and detailed information on specific topics. It is a measure of the model’s ability to delve into complexities, offer nuanced insights, and demonstrate expertise in narrow subject areas Bai et al. (2024). Thus, we can simply define a knowledge source $\mathcal{L}$ as $\mathcal{L}=(\mathcal{B},\mathcal{K})$.
While LLMs demonstrate striking knowledge breadth in extensive domains and areas, their knowledge depth for providing truly in-depth and expert-level output is still not promising Zhang et al. (2024); Bai et al. (2024). In this work, we propose to address this problem by first analyzing the relationship between knowledge breadth and depth, with prompts and responses in the alignment tuning dataset. Intuitively, a large and diverse set of instructions allows the LLMs to cover a wide range of topics and areas, thus expanding its knowledge breadth. Similarly, more responses of various quality provide LLMs with great opportunities to fully understand the question and explore the required knowledge, thus leading to LLMs being able to provide in-depth insights and responses.
However, if we take a closer look at alignment tuning datasets, there is a significant imbalance between the number of instructions and responses: a typical alignment tuning dataset is in the format of $\mathcal{D}=\{(x_{1},y_{1}^{1},y_{1}^{2}),...,(x_{n},y_{n}^{1},y_{n}^{2})\}$, in which each sample ($x,y^{1},y^{2}$) encompasses a prompt $x$, the winning response $y^{1}$ and the losing response $y^{2}$. In this dataset, the instruction number $n$ is usually in the tens of thousands, while the response number is simply 2, where we have $n>>2$. Here we argue that this imbalance between instruction and response numbers actually implies the unbalanced resource allocation for knowledge breadth and depth learning during alignment tuning time, limiting LLMs’ exploration for more in-depth knowledge.
2.2 Simple Balance
To validate our hypothesis in knowledge breadth and depth and further explore the impact of instruction and response number in LLM alignment tuning, we propose a simple balance method. This method involves first conducting knowledge breadth compression (KBC) by clustering and sampling a subset of instruction. After that, knowledge depth augmentation (KDA) is adopted to produce synthetic response pairs to uniformly augment knowledge depth. We provide more details in Section 3.1 and Section 3.2. Thus, we uniformly balance knowledge breadth and depth from $\mathcal{L}=(n,2)$ to $\mathcal{L}_{bal}=(n\times s,\frac{2}{s})$, where $s\in(0,1]$ is the scaling ratio for balancing.
Figure 3: Overview of BPO pipeline. BPO first selects representative prompts to reduce knowledge breadth through embedding-based clustering. Next, it generates responses using the SFT policy and employs GPT-4 to score these responses to uniformly construct response pairs. Subsequently, BPO samples response pairs to dynamically augment knowledge depth through gradient-based clustering. Finally, DPO is applied to the sampled data, which ensures efficient alignment.
2.3 Preliminary Experiment & Analysis
Preliminary experiment aims to investigate the results of compressing knowledge breadth $\mathcal{B}$ with various scaling ratios while correspondingly and uniformly increasing knowledge depth $\mathcal{K}$. We follow the experimental and evaluation settings in Section 4.1.
In Figure 2, we demonstrate that using embedding-based clustering to select only 1% to 10% of the representative instructions for alignment can achieve performance comparable to, and even surpassing, alignment on the full dataset while significantly reducing computational resource requirements. This finding is consistent with that of Zhou et al. (2024), indicating that data quantity is not the key factor. By focusing on samples at the centroids of clusters, we ensure the quality of the selected prompts while reducing the probability of redundant prompts being selected, thus achieving better performance with fewer prompts. This indicates that a small subset of carefully selected instructions could already satisfy the need of LLMs to expand the breadth of knowledge.
Furthermore, augmenting the knowledge depth of LLMs by constructing diverse responses for these selected prompts leads to further improvements in model performance. We also conduct experiments on HHRLHF dataset, as shown in Appendix A. These results highlight that balancing knowledge breadth and depth by dataset manipulation enables a more efficient and effective alignment of the model.
3 Our Method: Balanced Preference Optimization
In this section, we introduce our BPO method, which involves two core steps: Knowledge Breadth Compression (Section 3.1) and Dynamic Knowledge Depth Augmentation (Section 3.2).
3.1 Knowledge Breadth Compression
To reduce the breadth $\mathcal{B}$ of the knowledge, we aim to select the most representative subset of prompts $\mathcal{X_{\text{rep}}}\subset\mathcal{D}_{\text{org}}$. We follow the previous works and adopt a clustering method based on the prompts’ embedding vectors. Specifically, we first obtain the embedding vector $e_{i}$ using a pre-trained embedded $f$ and have $e_{i}=f(x_{i})$.
We then apply the K-means clustering algorithm to the set of embedding vectors $\{e_{1},e_{2},...,e_{n}\}$ to partition them into $C$ clusters $\{\mathcal{C}_{1},\mathcal{C}_{2},...,\mathcal{C}_{C}\}$. Within each cluster $\mathcal{C}_{c}$, we select the top $s$ fraction of prompts, where $s$ is the scaling ratio, based on Euclidean distance $d(e_{i},\mu_{c})$, whose embedding vectors are closest to the centroid $\mu_{c}$ and intuitively are the most representative:
$\mathcal{X}_{rep}=\bigcup_{c=1}^{C}\left\{x_{i}\in\mathcal{C}_{c}\ \bigg{|}\ \text{rank}\left(d(\mathbf{e}_{i},\boldsymbol{\mu}_{c})\right)\leq s\cdot|\mathcal{C}_{c}|\right\}$ (1)
3.2 Dynamic Knowledge Depth Augmentation
Knowledge Depth Augmentation
To augment the knowledge depth $\mathcal{K}$ for a given preference learning dataset, we first tune the base model with LoRA Hu et al. (2021) using a subset of seed data and obtian the supervised fine-tuned (SFT) policy $\pi_{\theta}$, more details are in Appendix B. Then, we produce multiple responses for each prompt $x_{i}\in\mathcal{X}_{\text{rep}}$ using the SFT policy $\pi_{\theta}$ and follow previous work to adopt the LLM-as-a-judge Zheng et al. (2023a); Gao et al. (2023); Li et al. (2024a) and reject sampling Khaki et al. (2024b) approach to construct high-quality response pairs. Furthermore, since the SFT policy has already been optimized for common scenarios, its responses are typically safe and exhibit minimal diversity, making it challenging to generate response pairs with significant score differences. To address this for the safety dataset, we employed a novel method by utilizing jailbreaking prompts to elicit more diverse responses. We analyze this method and its effects in Section 5.3.
Gradient Computation and Projection
Building upon the previously established uniform depth augmentation, we dynamically allocate depth for each prompt $x_{i}$ by utilizing the gradient features of each (prompt-response pair). Following the approach outlined by Xia et al. (2024), we employ Low-Rank Adaptation (LoRA) during the initial supervised fine-tuning (SFT) step to minimize the number of trainable parameters. Additionally, we implement random projection (Johnson, 1984) to reduce the dimensionality of the LoRA gradients. Specifically, the parameter update process, based on the Adam optimizer (Kingma, 2014), is as follows:
$\theta^{t+1}-\theta^{t}=-\eta_{t}\Gamma(z,\theta^{t})$ (2)
$\Gamma(z,\theta^{t})\triangleq\frac{m^{t+1}}{\sqrt{v^{t+1}}+\epsilon}$ (3)
$m_{t+1}=\beta_{1}m_{t}+(1-\beta_{1})\nabla\ell(z,\theta_{t})$ (4)
$v_{t+1}=\beta_{2}v_{t}+(1-\beta_{2})(\nabla\ell(z,\theta_{t}))^{2}$ (5)
where $\beta_{1},\beta_{2}$ are the hyperparameters, $\epsilon$ is a small constant, and $\Gamma(z,\theta^{t})$ represents the first-order expansion for the Adam dynamics. Subsequently, for a given data sample $z$ and model checkpoint $\theta^{t}$, we use random projection (Johnson, 1984) to map the high-dimensional LoRA gradients $\nabla\ell(z;\theta^{t})$ into a lower d-dimensional space, denoted as ^$\widehat{\nabla}\ell(z;\theta^{t})$. This accelerates the following clustering process and more details are provided in Appendix C. This gradient-based data representation allows us to measure the contribution of each response pair to the model’s optimization, enabling more efficient learning by dynamically adjusting the allocated knowledge depth resource for each prompt.
Dynamic Depth Allocation
Figure 4: Dynamic allocation of response pairs based on gradient clustering. Different prompts require varying numbers of pairs. While most prompts can be adequately represented with a single response pair, certain prompts demand a more comprehensive exploration involving additional pairs.
We perform K-means clustering on the set of gradient features ^$\widehat{\nabla}\ell(z;\theta_{t})$ into $G$ clusters $\{G_{1},G_{2},\ldots,G_{G}\}$ to evaluate the required knowledge depth for each sample from the model optimization perspective. For each cluster $G_{g}$, we select the top $\eta\%$ of (prompt, response pair) whose gradient vectors are closest to the cluster centroid $\boldsymbol{\mu}_{g}$, based on the Euclidean distance:
$\begin{split}\mathcal{D}_{dyn}=\bigcup_{g=1}^{G}\left\{z_{i}\in\mathcal{G}_{g}\ \bigg{|}\ \text{rank}\left(d(\widehat{\nabla}\ell(z;\theta^{t}),\boldsymbol{\mu}_{g})\right)\leq\eta\cdot|\mathcal{G}_{g}|\right\}\end{split}$ (6)
More results and analysis about the hyperparameter $\eta\%$ are shown in Table 5. Intuitively, informative augmented samples with gradient vectors close to the centers will get a greater opportunity to be selected, thus being allocated with more depth learning resources. Figure 4 shows the dynamic allocation of response pairs for different prompts, where the number of pairs varies based on the gradient-based clustering and selection. This illustrates how BPO dynamically adjusts the number of response pairs for different samples to optimize the knowledge depth allocation. We then perform DPO fine-tuning using $\mathcal{D}_{dyn}$, leading to better alignment and performance improvements as demonstrated in our experiments.
Model Method safeRLHF HH-RLHF
Data Size MT-Bench AlpacaEval Data Size MT-Bench AlpacaEval
[…]
KBC ($s$=10%) 5,094 5.12 45.75% 2,744 4.11 20.12%
Simple Balance 45,522 5.19 46.18% 32,928 4.42 25.34%
Simple Balance [1] 4,552 5.02 43.48% 3,292 4.21 23.75%
Curry-DPO 45,522 5.21 44.03% 32,928 4.67 28.29%
RS-DPO 52,067 5.32 47.48% 34,177 5.12 33.31%
BPO (Ours) 4,570 5.45 48.24% 3,312 5.48 40.25%
[…]
KBC ($s$=10%) 5,094 5.38 46.71% 2,744 5.14 22.85%
Simple Balance 29,555 5.68 50.30% 26,792 5.51 24.68%
Simple Balance [1] 2,955 4.02 48.64% 2,679 4.64 21.11%
Curry-DPO 29,555 5.67 48.50% 26,792 5.59 24.22%
RS-DPO 49,664 5.79 50.73% 27,143 5.83 25.26%
BPO (Ours) 3,042 5.85 54.27% 2,753 5.74 26.96%
Table 1: Performance comparison of different pair selection strategies on the SafeRLHF and HH-RLHF datasets. [1] stands for utilizing Simple Balance on 10% randomly selected samples. As shown, our method BPO can achieve comparable or even better performance with no more than 10% overall data.
4 Experiments
In this work, we introduce our experiment settings and analyze our results.
4.1 Experimental Setup
Datasets
We choose HH-RLHF Bai et al. (2022) and safeRLHF Dai et al. (2023) as our training sets. HH-RLHF consists of two subsets, with approximately 170K chosen-rejected pairs related to human and AI assistant dialogues. For our experiment, we randomly select 30,000 samples from the helpfulness subset. SafeRLHF contains 83.4K preference entries, where each entry includes two responses with labels for harmlessness and helpfulness. In our experiment, we select samples where both labels are the same, resulting in approximately 55,000 samples. Additionally, for each dataset, we randomly select 10% of the dataset $\mathcal{D}_{org}$ as the seed data $\mathcal{D}_{seed}$, using annotated chosen responses as golden reference to perform preliminary SFT.
Baselines
Two recent preference data optimization methods are used to compare with ours in this domain: RS-DPO (Khaki et al., 2024b) and Curry-DPO (Pattnaik et al., 2024b). RS-DPO initially generates diverse responses from a supervised fine-tuned model. Then it guides a reward model to identify contrastive pairs based on its reward distribution, and applies DPO to optimize the model’s alignment. In our experiment, we set the threshold of scoring variances as 0.9. Curry-DPO leverages multiple preference pairs per prompt, structured through curriculum learning from easy to hard tasks. It systematically ranks these pairs based on response quality differences and iteratively trains the model, resulting in improved performance across various benchmarks compared to standard DPO.
Implementation Details
We conduct experiments using Llama-2-13B (Touvron et al., 2023) and Llama-3-8B (Dubey et al., 2024). We also utilize Llama3-8b as the feature extractor to obtain representations used for clustering, which serves as a common choice in many clustering work (Petukhova et al., 2024). We adopt the MT-Bench (Zheng et al., 2023a) and AlpacaEval (Li et al., 2023) to evaluate the effectiveness of our proposed methods and baseline models. We set the scaling ratio for knowledge breadth and depth balance to 0.1. More evaluation details are attached in Appendix G and the training hyperparameters are shown in Appendix H. Additionally, we conducted further experiments on other datasets to validate the effectiveness and robustness of our method, which are detailed in Appendix I.
[…]
4.2 Experimental Results
In this section, we compare our overall BPO methods with the mentioned baselines and sampling methods. As shown in Table 1, BPO BPO achieves comparable or even superior performance while using no more than 10% of the data. KBC underperforms compared to BPO as it only relies on using fewer representative instructions, which further demonstrate the necessity for our hierichical sampling approach to balance between both breadth and depth. Compared to simple balance, BPO achieves better performance, which we attribute to the gradient-based dynamic knowledge depth augmentation selecting the most useful and informative samples for model updates. Additionally, potentially redundant samples are less likely to be selected during sampling. This reduces the possible negative impact of knowledge on the model and increases training efficiency. Furthermore, we compared our method with the approach of randomly selecting 10% of the samples used in simple balance. The results show that our gradient clustering performs better, indicating that BPO successfully allocates appropriate knowledge depth resources, thereby enhancing model performance. Furthermore, we compared our method with Curry-DPO and RS-DPO. Notably, RS-DPO’s performance ranks second only to ours, even surpassing BPO on MT-Bench when trained on the HH-RLHF dataset. However, it requires significantly more data than our approach. Overall, our method is both effective and efficient, demonstrating robustness across two datasets and two model sizes while successfully achieving alignment.
[…]
5 Extra Investigation & Further Analysis
In this section, we explore how different clustering configurations affect BPO’s performance, compare approaches for measuring knowledge depth, and analyze various response generation methods.
5.1 Clustering Settings
Number of Clusters
We investigate how the number of clusters affects both embedding vector clustering during the knowledge breadth compression step and gradient clustering in the dynamic knowledge depth augmentation step. Figure 5 shows that when using 100 clusters for embedding vector clustering and 50 clusters for gradient clustering, the results are relatively stable. A cluster number that is too small may fail to adequately represent the comprehensive level of knowledge breadth, making it difficult to achieve stable results and leading to significant fluctuations. Conversely, an excessively large number of clusters might over-segment the data, which does not necessarily guarantee a steady improvement in performance.
Clustering Algorithm
We explore the influence of different clustering algorithms on the performance of BPO. In addition to K-means, we apply K-medoids and spectral clustering. Figure 6 demonstrates that the performance difference between K-means and spectral clustering is minimal, while K-medoids performs slightly worse than the others. Furthermore, these clustering-based selection methods consistently outperform random selection, and except in rare cases, generally achieve better performance than using all data without selection. This further highlights the robustness of our BPO in various clustering algorithm selections.
5.2 Measurement of Required Knowledge Depth
In our BPO, we use a gradient-based method to estimate the required depth of knowledge for each sample. To ensure its efficacy and our method’s generalization, we also explore additional measurement approaches from various perspectives, including length and semantic similarity.
Length
To evaluate required knowledge depth, we measure the length of generated responses. We assume that high variability in response lengths for an instruction potentially may indicate inconsistent understanding, suggesting a need for deeper learning. Consequently, prompts with greater variance in response length are assigned greater depth. Further details are provided in Appendix E.
Semantic Similarity
We also utilize semantic similarity to measure required knowledge depth (Petukhova et al., 2024). We assume that high similarity in responses intuitively indicates a stable understanding, requiring less learning. Conversely, low similarity suggests inconsistent understanding, necessitating deeper learning. Further details are provided in Appendix F.
Measurement Approach Comparisons
Table 2 illustrates that relying solely on response length is inadequate, as it overlooks semantic differences. Responses of similar length can have different meanings, leading to an incomplete exploration of the data. Although semantic similarity performs better than response length, it still underperforms our gradient-based methods. While semantic similarity captures the correlations within the responses of each prompt, gradient feature represents knowledge breadth and depth on a global scale, resulting in more effective optimization.
Measurement SafeRLHF HH-RLHF
MT-Bench AlpacaEval MT-Bench AlpacaEval
Length 4.48 34.85% 4.27 24.61%
Semantic Similarity 4.92 40.87% 4.68 29.75%
Gradient 5.45 48.24% 5.48 40.25%
Table 2: Performance comparison of different methods for measuring the required knowledge depth. The results demonstrate that measuring via gradients in our BPO leads to better performance.
5.3 Comparisons about Augmentation Methods in KDA
As described in Table 3, common response generation methods adopted in many online DPO variants on the safeRLHF dataset surprisingly did not yield promising results. Upon analyzing the bad cases, we found that the base model had already been optimized for common scenarios, making its responses quite safe and difficult to obtain unsafe ones.
Therefore, we employed a different method in the safety alignment area: for half of the generations, we used jailbreaking prompts proposed in Shen et al. (2024) to induce the model to produce potentially risky responses. These were then paired with normally generated safe responses for further processing. The pair construction process is the same as described in Section 3.2. Our findings also support the assumption that uniformly increasing the number of more distinctly different pairs between chosen and rejected responses can notably enhance the model alignment effects. Moreover, in the context of online RLHF research focused on safety alignment, we want to emphasize to future research the importance of assessing quality differences in generations. It is crucial to determine whether the model is already sufficiently safe, as this could hinder the generation of adequately poor negative responses.
Generation Methods MT-Bench AlpacaEval
w/o data generation 5.12 45.75%
normal generations 4.40 31.18%
half normal / half jailbreak 5.19 46.18%
[…]
6 Related Work
Preference Learning. Preference Learning aims to align LLMs’ output distribution with human preferences and values. The most widely used alignment method is Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017), which involves training a reward model based on human feedback. This reward model is then utilized within a Proximal Policy Optimization (PPO) framework (Schulman et al., 2017) to fine-tune LLMs, ensuring their outputs align more closely with human preferences. However, this complex pipeline often suffers from instability and poor sample efficiency (Zhu et al., 2023b) To address these issues, several enhanced offline approaches have been proposed, such as SLiC (Zhao et al., 2022), SLiC-HF (Zhao et al., 2023a) and so on. Direct preference optimization (DPO) (Rafailov et al., 2024) directly optimizes the language model by directly training from the policy differences between chosen and rejected pairs. More simple and direct alignment variants are proposed afterwards (Meng et al., 2024; Ethayarajh et al., 2024). Specifically, curry-DPO (Pattnaik et al., 2024c) suggests that using diverse featured data can significantly impact alignment performance. By integrating curriculum learning with DPO, it effectively enhances performance.
Preference Data Optimization. The preference data often encompasses a prompt, a chosen response, and a rejected response Tan et al. (2024). As underscored in (Morimura et al., 2024), typical offline alignment methods are highly sensitive to the quality and coverage of preference data (Kalai and Vempala, 2024). Moreover, selecting a dataset with the right balance of size, complexity, quality, and diversity can significantly impact alignment effectiveness (Liu et al., 2023; Zhou et al., 2024; Zhao et al., 2023b; Wei et al., 2023; Song et al., 2024a; Xiong et al., 2024). Hence, more representative user queries and corresponding optimized preference pairs might be more helpful and lead to less propagated problems (Song et al., 2024b). In this way, many previous work utilize a reward model to rank and select preference pairs (Cui et al., 2024; Zhu et al., 2023a; Liu et al., 2023; Morimura et al., 2024) Notably, Liu et al. (2024) statistically proves utilizing rejection sampling to select preference pairs can greatly benefit to a better alignment performance. Khaki et al. (2024b); Touvron et al. (2023) systematically combine rejection sampling when constructing pairs, which shows remarkable improvements of both performance and data efficiency.
7 Conclusions
In this work, we expose the potential imbalance between knowledge breadth and depth learning within preference learning datasets, and we conduct preliminary experiments using a simple balancing method to validate this observation. Furthermore, we propose a novel Balance Preference Optimization (BPO) method, which dynamically enhances the knowledge depth learning resources for each sample. BPO measures the knowledge depth of each sample from a unique model optimization perspective and dynamically selects varying numbers of response pairs for each preference sample. Our experimental results on several alignment benchmarks demonstrate that BPO outperforms other preference optimization methods while maintaining high training efficiency. Additionally, we provide an in-depth analysis of BPO, offering insights and guidelines on balancing preference learning datasets for future research.
Limitations and Future Work
Despite the promising results of our hierarchical clustering-based framework, several limitations must be acknowledged. First, learning directly from the model’s generated responses leads to limited or marginal performance improvement, indicating inherent constraints in self-optimization. To achieve high-quality data curation, we rely on the GPT-4 model to score responses to construct response pairs, which introduces a certain level of dependence on the performance and potential biases of GPT-4.
For future work, it is crucial to develop statistical or empirical methods to evaluate whether a user query would be beneficial to the entire alignment pipeline. Additionally, refining the generation and construction of preference pairs is essential, as it fundamentally shapes the alignment learning space. Further exploration into these areas could yield significant advancements in the alignment capabilities of LLMs.
Appendix A Additional Results for Preliminary Experiment
Scaling Ratio safeRLHF HH-RLHF
MT-Bench (Avg Score) AlpacaEval (win %) MT-Bench (Avg Score) AlpacaEval (win %)
Fullset 4.90 36.84% 4.71 18.14%
[…]
1% Simple Balance 5.17 41.22% 4.29 21.41%
[…]
5% Simple Balance 5.03 42.43% 4.34 24.61%
[…]
10% Simple Balance 5.19 46.18% 4.42 25.34%
[…]
20% Simple Balance 5.10 45.47% 4.36 22.30%
[…]
Appendix B Preliminary Supervised Fine-Tuning
The preliminary step involves fine-tuning the unsupervised base model $\pi_{0}$ using a small, high-quality annotated seed data $\mathcal{D}_{seed}=\{(x_{1},y_{1}),\dots,(x_{m},y_{m})\}$, where $x_{i}$ is the input text and $y_{i}$ is the corresponding target output. Additionally, we use LoRA, which freezes the pre-trained model weights and injects trainable rank decomposition matrices into each layer, to reduce parameters that need to be trained (Hu et al., 2021). This supervised fine-tuning (SFT) step effectively refines the base model, enhancing the subsequent DPO performance (Wang et al., 2024; Saeidi et al., 2024). The fine-tuning objective is to minimize the supervised loss:
$\mathcal{L}_{\text{SFT}}(\theta)=\mathbb{E}_{(x_{i},y_{i})\sim\mathcal{D}_{\text{seed}}}\left[-\log\pi_{\theta}(y_{i}\mid x_{i})\right]$ (7)
where $\theta$ represents the model parameters.
Appendix C Random Project for Gradients
Following Xia et al. (2024), for a given data sample $z$ and model $\theta$, we compute a $d$-dimensional projection of the LoRA gradient ^$\widehat{\nabla}\ell(z;\theta_{t})=\Pi^{\top}\nabla\ell(z;\theta_{t})$, where each entry of $\Pi\in\mathbb{R}^{P\times d}$ is drawn from a Rademacher distribution (Johnson, 1984) (i.e., $\Pi_{ij}\sim\mathcal{U}(\{-1,1\})$). In total, we compute the projected gradient features for each data sample $z$ as $\widetilde{\Gamma}(z,\cdot)=\Pi^{\top}\Gamma(z,\theta_{t})$.
Appendix D Additional Results for Dynamic Knowledge Depth Augmentation
Number of Clusters $\eta\%$ MT-Bench AlpacaEval
50 1% 5.19 26.46%
5% 5.32 33.79%
10% 5.48 40.25%
100 1% 4.30 25.59%
5% 4.67 32.17%
10% 5.29 34.16%
Table 5: Influences of different $\eta\%$ to dynamic knowledge depth augmentation stage in our method, BPO. Clustering method is K-means, and the number of clusters is 50 and 100. Experiments were conducted on Llama-3-8B, HH-RLHF dataset.
Appendix E Length Differences
For generated responses, we consider their token length to measure the required knowledge depth. For each prompt $x_{i}$, let $\{y_{i1},y_{i2},\dots,y_{ik}\}$ denote the generated candidate responses. We compute the token length $l_{ij}$ of each response $y_{ij}$. The sample variance of the lengths for prompt $x_{i}$ is calculated as:
$\sigma^{2}_{l_{i}}=\frac{1}{k-1}\sum_{j=1}^{k}(l_{ij}-\bar{l}_{i})^{2}$ (8)
where $\bar{l}_{i}=\frac{1}{k}\sum_{j=1}^{k}l_{ij}$ is the mean length of the responses for prompt $x_{i}$. This variance $\sigma^{2}_{l_{i}}$ reflects the diversity in response lengths, indicating the consistency level of understanding. These variances are normalized to obtain the weight $w_{i}^{l}$ for each prompt:
$w_{i}^{{l}}=\frac{\sigma^{2}_{l_{i}}}{\sum_{i=1}^{n^{{}^{\prime}}}\sigma^{2}_{l_{i}}}$ (9)
[…]
Appendix F Semantic Similarity Difference
We also leverage semantic similarity to measure the required knowledge depth. For each prompt $x_{i}$, let $\{y_{i1},y_{i2},\dots,y_{ik}\}$ denote the set of generated candidate responses. We obtain the embedding vector $\mathbf{e}_{ij}$ for each response $y_{ij}$ using a pre-trained embedder $f$. We compute the pairwise cosine similarities between all response embeddings and calculate the average cosine similarity $\bar{s}_{i}$ for prompt $x_{i}$:
$\bar{s}_{i}=\frac{2}{k(k-1)}\sum_{1\leq p<q\leq k}\cos\left(\mathbf{e}_{ip},\mathbf{e}_{iq}\right)$ (10)
where
$\cos\left(\mathbf{e}_{ip},\mathbf{e}_{iq}\right)=\frac{\mathbf{e}_{ip}\cdot\mathbf{e}_{iq}}{\left\|\mathbf{e}_{ip}\right\|\left\|\mathbf{e}_{iq}\right\|}$ (11)
[…]
$w_{i}^{s}=\frac{\dfrac{1}{\bar{s}_{i}}}{\sum_{i=1}^{n^{\prime}}\dfrac{1}{\bar{s}_{i}}}$ (12)
where $n^{\prime}$ is the number of prompts in $\mathcal{X}_{ep}$. This inverse relationship ensures that prompts with higher overall response similarity (i.e., less semantic diversity) are assigned a smaller depth. Based on these weights, we allocate the depth $k_{i}$ for each prompt $x_{i}$ as:
$k_{i}=\left\lceil w_{i}^{s}\times N\right\rceil$ (13)
[…]
Appendix G Dataset Details
Trainset
HH-RLHF consists of two subsets, with approximately 170K chosen-rejected pairs related to human and AI assistant dialogues. We randomly select 30,000 samples from the helpfulness subset for our experiment. SafeRLHF involves 83.4K preference entries and each is designed with two responses, accompanied by two labels for harmlessness and helpfulness. In our experiment, we select samples where the two labels are the same, yielding approximately 55,000 samples. Finally, for each dataset, we randomly select 10% of the dataset $\mathcal{D}_{org}$ as the seed data $\mathcal{D}_{seed}$, using annotated chosen responses as golden reference to perform preliminary SFT.
Evaluation Set
Specifically, MT-Bench is a multi-turn benchmark that evaluates the capacity of LLMs to engage in coherent, informative, and engaging conversations. We present the average score of each round. AlpacaEval, on the other hand, is designed to evaluate the win-lose-tie rate of LLMs by comparing their responses to the baseline model, text-davinci-003.
Appendix H Training Hyperparameters
The learning rate is set to 2e-5, with a batch size of 32 for 4 epochs. In the LoRA setting, we use the AdamW optimizer. When constructing pairs, we follow Khaki et al. (2024a) to generate responses for 10% of representative prompts sampled through embedding-based clustering as described in Section 3.2. We set k = 16 responses per prompt, with a maximum of 512 new tokens, a top-k value of 50, and a temperature of 1. Then, we select the top 10% of pairs with the highest score differences from all constructed pairs, and compute the gradients of the selected samples. For gradient projection, we follow Xia et al. (2024)’s settings to randomly project the gradient features to d = 8192. In DPO training, we set the learning rate to 2e-5, with a batch size of 16 for 4 epochs.
Appendix I Additional Experiments
To further demonstrate that BPO is both effective and robust, we conducteded additional experiments on the UltraFeedback dataset Cui et al. (2023), anselectedcted AlpacaEval 2.0 Dubois et al. (2024) and Arena-Hard Li et al. (2024b) as our evaluation benchmarks. All experiments were carried out using the Llama-3-8B-Instruct model.
Trainset
UltraFeedback contains AI-generated feedback annotations for a wide variety of prompts, facilitating preference modeling. For our experiments, we randomly select 30,000 samples from UltraFeedback, of which 10% are used as seed data $\mathcal{D}_{seed}$, following the setup described in Section 4.1. We used the same training hyperparameters as presented in the Appendix H.
Evaluation Set
AlpacaEval 2.0 comprises 805 questions drawn from five different datasets, while Arena-Hard includes 500 well-defined technical problem-solving queries. We report all scores according to the respective evaluation protocols of the benchmarks. For AlpacaEval 2.0, we report both the raw win rate and the length-controlled win rate (LC), where the LC metric is designed to be robust against model verbosity. For Arena-Hard, we report the win rate compared to the baseline model.
Method AlpacaEval 2.0 Arena Hard
LC Win Rate (%) Win Rate (%) Win Rate (%)
Base-model 22.92 22.57 20.20
Vanilla DPO 28.71 27.24 33.21
Curry-DPO 35.46 36.65 34.71
RS-DPO 41.23 40.44 37.98
BPO (ours) 44.75 43.68 39.83
Table 6: Performance comparison of different pair selection strategies on the UltraFeedback dataset.
As shown in Table 6, our BPO method demonstrates substantial improvements over the base model and other baselines on both the AlpacaEval 2.0 and Arena Hard datasets. These additional experiments further validate the robustness and effectiveness of BPO across diverse datasets and models.
Appendix J LLM-as-a-Judge’s Prompt
We follow the prompt designed in (Zheng et al., 2023b). Prompts used in AlpacaEval are shown as follow:
{mdframed}
[System]
Please act as an impartial judge and evaluate the quality of the responses
provided by two AI assistants to the user question displayed below.
You should choose the assistant that follows the users instructions and answers
the users question better. Your evaluation should consider factors
such as the helpfulness, relevance, accuracy, depth, creativity, and level of
detail of their responses. Begin your evaluation by comparing
the two responses and provide a short explanation. Avoid any position biases
and ensure that the order in which the responses were presented does not
influence your decision. Do not allow the length of the responses to influence
your evaluation. Do not favor certain names of the assistants. Be as objective
as possible. After providing your explanation, output your final verdict by
strictly following this format: "[[A]]" if assistant A is better,
"[[B]]" if assistant B is better, and "[[C]]" for a tie.
[User Question]  {}
[The Start of Assistant As Answer] {} [The End of Assistant As Answer]
[The Start of Assistant Bs Answer] {} [The End of Assistant Bs Answer]
Prompts designed to evaluate models’ performance in MT-Bench are attached below:
{mdframed}
[System] Please act as an impartial judge and evaluate the quality
of the response provided by an AI assistant to the user question displayed below.
Your evaluation should consider factors such as the helpfulness, relevance,
accuracy, depth, creativity, and level of detail of the response.
Begin your evaluation by providing a short explanation.
Be as objective as possible. After providing your explanation,
please rate the response on a scale of 1 to 10 by strictly following this format:
"[[rating]]", for example: "Rating: [[5]]".
[Question] {}
[The Start of Assistant’s Answer] {}
[The End of Assistant’s Answer]
Appendix K Acknowledgment of AI Assistance in Writing and Revision
We utilized ChatGPT-4 for revising and enhancing wording of this paper.